Given the function h described by h(x)=x+5 find each of the following
h(0)
h(-8)
h(-3)
h(7)
h(c+14)

Answer by Alan3354(69443)

Substitute the values for x.
h(x) = x+5
h(0) = 0+5 = 5
h(-8) = -8+5 = -3
h(-3) = -3+5 = 2
h(7) = 7+5 = 12
h(c+14) = c+14+5 = c+19